OpenStudy (anonymous):

Im doing substitution right now in calcalus and im kinda confused on how to do this problem. {integral}sqrt(x)sin^2(x^(3/2)-1)dx. u=x^(3/2)-1 Im doing substitution right now in calcalus and im kinda confused on how to do this problem. {integral}sqrt(x)sin^2(x^(3/2)-1)dx. u=x^(3/2)-1 @Mathematics

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{x}\sin^2(x^{3/2}-1)dx\] \[u=x^{3/2}-1\] \[du=(3/2)\sqrt{x}\] \[(3/2)\int\limits_{}^{}\sin^2(u)du\] Evaluate this integral and substitute \[u = x^{3/2}\]

OpenStudy (anonymous):

so how did you get rid of the sqrt(x)?

OpenStudy (anonymous):

It "disappears" because it is part of du.

OpenStudy (anonymous):

ahhh ok i see it tyvm